OpenStudy (anonymous):

the function D(x) models the cumulatie number of deaths from a diease x years after 1984. Estimate the year when there were 88,000 deaths

OpenStudy (anonymous):

are you going to tell us what the function is?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

you need to tell us what the function D(x) is if you want us to estimate something with it.

OpenStudy (anonymous):

the function

OpenStudy (anonymous):

the function D(x) models the cumlative number of deaths , this is all the question says...

OpenStudy (anonymous):

then how is it possible to estimate the year there were 88000 deaths?!

OpenStudy (anonymous):

there must be a definition of D(x). look around.

OpenStudy (anonymous):

D(x)= 3332x^2+5060x+5855 sorry I am confused, i have been working on this all day...

OpenStudy (anonymous):

OK, so you need to find the first year D(x) exceeds 88000, yes?\[3332x^2 + 5060x + 5855 = 88000\]\[3332x^2 + 5060x - 74145 = 0\]Use the quadratic formula to solve for x.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or do you want the number of deaths in a particular year, not the cumulative number that year?

OpenStudy (anonymous):

ti appreciate your patience, but the question ask me to estimate the year when there were 88,000

OpenStudy (anonymous):

right. so then take the derivative to get rate of change and get 6664x + 5060. set that equal to 88000 and get: 6664x + 5060 = 88000 6664x = 82940 x = 12.44